For its inspecting cost pool, Ellsworth, Inc. expected overhead cost of $400,000 and 4,000 inspections. The actual overhead cost for that cost pool was $480,000 for 5,000 inspections. The activity-based overhead rate used to assign the costs of the inspecting cost pool to products is

$120 per inspection.
$80 per inspection.
$96 per inspection.
$100 per inspection.
Answer #1
Concepts and reason

Cost: Cost is any value spent to produce a product or to render any service. The types of costs are fixed cost and variable cost.

Overhead cost: Overhead costs are the indirect cost incurred to operate the business. They are not connected with the production of goods or rendering any service. They do not include any direct cost.

Fundamentals

Activity based costing: Activity based costing is a method of costing that assigns cost to products. Indirect costs are recognised through it. This is mostly used by manufacturing industries. The cost of products that has huge materials involved are transferred to products of less contents. Thus it raises the low volume cost unit.

Cost pool: Factory overheads are allocated to production units. It is the process of sorting into groups. Individual costs are merged as groups. It is used in activity-based costing.

Factory overhead: Factory overhead costs are the indirect cost incurred to manufacture a product. It is also known as work cost. It is classified into variable factory overheads and fixed factory overheads. Variable factory overheads only include the variable cost and fixed factory overhead include the fixed costs.

Justification of incorrect answers:

The product of activity based overhead rate and number of inspections should be equal to expected overhead costs.

On multiplying $120, $80 and $96 the product is not $400,000 of overhead costs.

Therefore, these answers are incorrect.

Calculate the activity based overhead rate:

Activitybasedoverheadrate}=ExpectedoverheadcostExpectedinspections=$400,0004,000=$100perinspections\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Activity based }}\\\\{\rm{overhead rate}}\\\end{array} \right\} = \frac{{{\rm{Expected overhead cost}}}}{{{\rm{Expected inspections}}}}\\\\ = \frac{{\$ 400,000}}{{4,000}}\\\\ = \$ 100{\rm{ per inspections}}\\\end{array}

Therefore, activity based overhead rate is $100 per inspection.

Ans:

Activity based overhead rate is $100 per inspection.